OpenStudy (cannah):

A charge +q is located at the origin, while an identical charge is located on the x axis at x = +0.83 m. A third charge of +2q is located on the x axis at such a place that the net electrostatic force on the charge at the origin doubles, its direction remaining unchanged. Where should the third charge be located?

OpenStudy (skylyerbailey):

The answer might be 1.0716m When+q is placed at the origin and an identical charge is placed 0.76 m from it.....the net force on the charge placed at origin= F1 = kq^2/(0.76)^2 to double the force on the charge at the origin charge +2q must be placed on the x axis such that F2(due to +2q) + F1 = 2F1 the distance between +2q and +q(at origin) is x NEXT.....k*2q*q/x^2 = k*q^2/(0.76)^2 then after simplifying x^2 = 2(0.76)^2.... square root both sides x= sqrt(2) * 0.76 =1.41*0.76 = 1.0716

OpenStudy (skylyerbailey):

I hope this helps! :)